question 1
The slope of a line is 12 and the point (1, 10) is on the line. Express the line in slope-intercept form.
question 2
The slope of a line is 11 and the point (14, 19) is on the line. Express the line in slope-intercept form.

Found 2 solutions by CharlesG2, solver91311:
Answer by CharlesG2(834)

question 1
The slope of a line is 12 and the point (1, 10) is on the line. Express the line in slope-intercept form.
question 2
The slope of a line is 11 and the point (14, 19) is on the line. Express the line in slope-intercept form.

slope-intercept form is y = mx + b, where m is the slope and b is the y-intercept (vertical-intercept or point (0,b))

1.) y = mx + b, plug in m = 12, x = 1, y = 10, solve for b
10 = 12(1) + b
10 = 12 + b
10 - 12 = b
-2 = b
the equation of the line is y = 12x - 2

2.) y = mx + b, plug in m = 11, x = 14, y = 19, solve for b
19 = 11(14) + b
19 = 154 + b
19 - 154 = b
-135 = b
the equation of the line is y = 11x - 135

Answer by solver91311(24713)


One question per post. But since these are really just the same question with different numbers, I'll show you how to answer any question like this.

The slope intercept form is

where is the slope and is the -intercept, that is is the value of the -coordinate of the ordered pair that represents the point where the graph of the line intersects the -axis.

So, to solve either of your questions, first substitute the given slope number for . Then take the -coordinate of the given point and substitute for in . And then take the -coordinate of the given point and substitute for in .

Making those substitutions will leave you with an equation that has as a variable. Solve for and you will have sufficient information to write your slope intercept equation.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism